OpenStudy (kaylala):

topic: PROVING IDENTITIES -- new -- (SEE COMMENTS) (trigonometry) i'm practicing how to solve :)

OpenStudy (kaylala):

1) cot A + 2 cos A = [(2 sin A + 1)/(tan A)] 2) [(2+csc x)/(sec x)] - 2 csc x = cot x 3) (1)/(sec x - tan x) = sec x + tan x

OpenStudy (ranga):

cot A + 2 cos A = 1/tan(A) + 2cos(A) * sin(A)/sin(A) = 1/tan(A) + 2sin(A)/tan(A) = (1 + 2sin(A))/tan(A)

OpenStudy (kaylala):

@ranga how about #2 and #3?